OpenStudy (jolo_yolo):

The present cost of a van is $9,000. The cost of this van depreciates at the rate of $800 per year. Part A: Write a function to show the cost of the van f(t) after t years. Part B: What is the total cost of the van after 3 years? Part C: If the cost of the van was $500 less than the present cost, what would be the cost of the van after 6 years?

OpenStudy (ccswims):

for part A : If it goes down the slope will be a negative, do you know the slope?

OpenStudy (jolo_yolo):

I did my whole test and now im stuck with this and another question ._.

OpenStudy (ccswims):

are you in flvs?

OpenStudy (ccswims):

i remember this test

OpenStudy (ccswims):

Ok, Part A: f(t) = -800t + 9,000 will be our function

OpenStudy (ccswims):

For Part B: Just plug in 3 for "t" to get the answer... f(3)= -800(3) + 9,000 Can you do this by yourself?

OpenStudy (jolo_yolo):

Yes, I can

OpenStudy (ccswims):

Ok, and Part C: Just subtract 500 to 9,000 9,000 - 500= 8,500 Now just do what we did in Part B except replace 9,000 with 8,500, so f(6) = -800(6) + 8,500

OpenStudy (jolo_yolo):

So for Part B it's f(3) = 11400 and for Part C it's f(6) = 13300 - right?

OpenStudy (jolo_yolo):

@ccswims

OpenStudy (ccswims):

No, part B will be 6,600

OpenStudy (ccswims):

I think you multiplied 800 as a positive, it should be a negative, -800

OpenStudy (ccswims):

and Part C will be 3,700

OpenStudy (jolo_yolo):

Oops, yea I did. Sorry, thanks for the help! Can you help wit one more?

OpenStudy (ccswims):

yea sure

OpenStudy (ccswims):

sorry I took awhile, I'm taking a test of my own and I'm jumping back and forth lol

OpenStudy (jolo_yolo):

lol it's fine, Im quite patient. Part A: Mary rented a stage for 4 days and it costs her $410. If she would have rented it for 8 days, it would have cost $770. Write an equation in the standard form to represent the total rent (y) that Mary has to pay for renting the stage for x days. Part B: Write the equation obtained in Part A using function notation. Part C: How would you graph this function? Please include information on your labels, axes and intervals.

OpenStudy (ccswims):

Part A: Standard Form = Ax + By = C so we know our equation is y = 90x + 50, now we must switch it to standard form

OpenStudy (ccswims):